OpenStudy (anonymous):

A red blood cell has a volume of 150 micrometer cubed. A 70 kg person has approx. 5.0 litres of blood, 45 percent of which is red blood cells. How many red blood cells are in the body?

OpenStudy (anonymous):

what I did first was find how much RBC are in 5.0 litres of blood...... then I jhave no idea lol

OpenStudy (jebonna):

Erm this is just a guess but i worked out how many litres of RBC there are in blood which was 2.25 litres and then do you convert that number into micrometers to get the answer? Im not 100% lol good luck though

OpenStudy (aaronq):

First we find the volume of red cells in 5 L of blood. 0.45*5 L=2.25 L then we divide the total volume of RBC by the volume of 1 RBC, but first we have to convert to the same units. using the conversion factor \(\sf 1 ~\mu m=1.0*10^{-5} d m\) \(\large V_{RBC} = \sf 2.25~dm^3*(\frac{1~\mu m}{1.0*10^{-5}dm})^3=2.25*10^{24}\mu m^3\) \(\sf number ~of~RBC=\dfrac{V_{RBC}}{V_{one~cell}}=\dfrac{2.25*10^{24}\mu m^3}{150~\mu m^3/cell }=1.5*10^{22}cells\)